# Analogs of the lower and upper central series in skew braces: a survey

### v3

@@ 1 Introduction
$a\circ(b\cdot c)=(a\circ b)\cdot a^{-1}\cdot(a\circ c)$ (1.1)
@@ 4 Analogs of the upper central series
Since $\lambda$ is a homomorphism on $(A,\circ)$, its kernel $\ker(\lambda)$ is a subgroup of $(A,\circ)$. But $\cdot$ and $\circ$ coincide on $\ker(\lambda)$. It then follows that $\ker(\lambda)$, and in particular $\mathrm{Soc}(A)$ in view of (4.1), is a subgroup of $(A,\cdot)$.
@@
    $\mathrm{Soc}(A)$ is normal in $(A,\circ)$: This follows from (4.3).
@@
We know that $\mathrm{Soc}(A)$ is a subgroup of $(A,\circ)$ by Proposition 4.1, so clearly $\mathrm{Ann}(A)$ is also a subgroup of $(A,\circ)$ by (4.2). Since $\cdot$ and $\circ$ coincide on $\ker(\lambda)$, it follows that $\mathrm{Ann}(A)$ is a subgroup of $(A,\cdot)$.
@@
$\mathrm{Soc}_{0}(A)=1,\quad\mathrm{Soc}(A/\mathrm{Soc}_{n}(A))=\mathrm{Soc}_{n+1}(A)/\mathrm{Soc}_{n}(A),$ (4.4)
@@
$\mathrm{Ann}_{0}(A)=1,\quad\mathrm{Ann}(A/\mathrm{Ann}_{n}(A))=\mathrm{Ann}_{n+1}(A)/\mathrm{Ann}_{n}(A).$ (4.5)
@@
$A_{1}=A,\quad A_{n+1}=A_{n}/\mathrm{Soc}(A_{n})$ (4.6)
instead, and the SocleSeries command in GAP [26] returns this series (4.6) rather than the series (4.4). But observe that by induction, we have a skew brace isomorphism, i.e. a map that is an automorphism with respect to both of the operations $\cdot$ and $\circ$, between
@@
Hence, the terms in the series (4.6) are nothing but quotients of those in the socle series of our definition, and there is no fundamental difference between the two series.
@@
$\mathrm{Ann}_{n}(A)\subseteq\mathrm{Soc}_{n}(A)$ (4.7)
@@
Let $A=(C_{p}\times C_{q},\cdot,\circ)$ be the brace in Example 3.1(i). From (4.1), we see that $\mathrm{Soc}(A)=\ker(\lambda)$ because $(A,\cdot)$ is abelian. Moreover, from (3.1), it is easy to see that $\mathrm{Soc}(A)=C_{p}\times\{1\}$ and the quotient $A/\mathrm{Soc}(A)$ is a trivial brace. This means that
@@
$\mathrm{Ann}_{n}(A)\subseteq\mathrm{Soc}_{n}(A)\subseteq\zeta_{n}(A,\cdot)\quad\mbox{and}\quad\mathrm{Ann}_{n}(A)\subseteq\zeta_{n}(A,\circ),$ (4.8)
@@ 5 Comparison of the different types of nilpotency
Nevertheless, there are some relations among the four different types of nilpotency. For example, from (4.7), it is clear that annihilator nilpotent implies socle nilpotent. Moreover, it was recently proven in [15] that a left nilpotent skew brace $A$ with nilpotent additive group $(A,\cdot)$ and $A^{3}=1$ is necessarily right nilpotent. But as we can see from Example 3.1, in general, there is no implication between left nilpotent and right nilpotent. Here, we shall survey a couple of known results about the various types of nilpotency.
@@ 2. $(b)$
    by (5.1), so it follows that $x\in\mathrm{Soc}_{c+j+1}(A)$, as desired.
@@
Now, assume that (b) holds. We have $\mathrm{Soc}_{n}(A)=A$ for some $n$ from the hypothesis. Note that $(A,\cdot)$ is nilpotent by (4.8). We also have
@@ 3. $(c)$
Here we assume that $A$ is left nilpotent. Then $Q$ is also left nilpotent, so by induction, we can assume that $(Q,\circ)$ is nilpotent. This means that
@@
because $\mathrm{Soc}(A)$ is normal in $(A,\circ)$ by Proposition 4.1. On the other hand, recall from (4.3) that $a\circ x\circ\overline{a}=\lambda_{a}(x)$ and this lies in $\ker(\lambda)$. This yields
$[a,x]_{\circ}$ $=\lambda_{a}(x)\circ\overline{x}$
$=\lambda_{a}(x)x^{-1}$
$=a*x.$
@@
Here we assume that $(A,\circ)$ is nilpotent. Then $(Q,\circ)$ is also nilpotent, so by induction, we can assume that $Q$ is annihilator nilpotent. This means that
@@
    by (5.2), which means that $x\in\mathrm{Ann}_{c+j+1}(A)$, as desired.
@@
Finally, we assume that $A$ is annihilator nilpotent, that is $\mathrm{Ann}_{n}(A)=A$ for some $n$. Note that $(A,\cdot)$ is nilpotent by (4.8). We also have
@@ 6 Relationship among the terms in the series
$a*y\in\mathrm{Ann}_{n}(A)*A^{n-(k+1)}\subseteq\mathrm{Ann}_{k+1}(A)$ (6.1)
@@
$=\varphi_{a}(\lambda_{x}(y))\varphi_{a}(y)^{-1}$ (6.2)
@@
and together with (6.1), this yields
@@
which can be assumed to be a homomorphism by induction on $k$ as in (6.1). In order to prove (F), we would have to show that
@@
as in (6.2) because $\lambda_{y}(x)$ and $x$ are not always elements of $A^{(n-(k+1))}$ here. There is no way to amend this problem. In fact, in general, (F) is false.
@@
$\mathrm{Im}(\psi_{b}-\mathrm{id}_{C})\subseteq\ker(\phi),$ (6.3)
@@
$(b,c)*(x,y)$ $=((\phi_{-c}-\mathrm{id}_{B})(x),(\psi_{b}-\mathrm{id}_{C})(y))$ (6.4)
$[(b,c),(x,y)]$ $=((\mathrm{id}_{B}-\phi_{y})(b)+(\phi_{c}-\mathrm{id}_{B})(x),0)$ (6.5)
@@
This proves the identities (6.4) and (6.5).
@@
with equalities when $b\in\langle\vec{e}_{3}\rangle$ but $b\neq\vec{0}$. We see that (6.3) is satisfied.
On the one hand, it is clear from (6.4) and the above that
@@
On the other hand, it is not hard to use (6.4) and (6.5) to verify that
$\mathrm{Ann}_{1}(A)$ $\supseteq\langle\vec{e}_{1}\rangle\times\langle\vec{e}_{1}\rangle,$ (6.6)
@@
The three inclusions in (6.6) follow from these observations.
Now, a simple calculation using (6.4) yields
@@ 7 Another lower central series and annihilator nilpotency
$\Gamma_{1}(A)=A,\quad\Gamma_{n+1}(A)=\langle\Gamma_{n}(A)*A,\,A*\Gamma_{n}(A),\,[A,\Gamma_{n}(A)]\rangle,$ (7.1)
@@
Therefore, perhaps (7.1) and the annihilator series (4.5) are, respectively, the “correct” analogs of the lower and upper central series, in the setting of skew braces. There is another reason to believe that this is indeed the case. A categorical approach was employed to study skew braces in [18], where it was shown that skew braces form a semi-abelian category, and the so-called Huq commutator of ideals is given as follows. Let us remark in passing that the categorical aspects of Hopf braces, an algebraic structure defined in [2] as a Hopf-theoretic generalization of skew braces, were also studied in [27].
@@
$\Gamma_{1}^{\prime}(A)=A,\quad\Gamma_{n+1}^{\prime}(A)=[A,\Gamma_{n}^{\prime}(A)]^{A},$ (7.2)
for all $n\in\mathbb{N}$. This coincides with the series (7.1). We note that Remark 7.6 and Proposition 7.7 were observed in [7] as well.
@@
Therefore, the series (7.1) can be defined using a commutator operator in the same way the lower central series of a group is defined. Our discussion so far also suggests that annihilator nilpotency is perhaps the more appropriate analog of nilpotency in the setting of skew braces.
@@
Let $I$ be an ideal of $A$. According to (7.2) and Proposition 7.7, as a skew brace, the lower central series of $I$ is defined by

### v4

@@ 1 Introduction
$a\circ(b\cdot c)=(a\circ b)\cdot a^{-1}\cdot(a\circ c)$ (1)
@@ 4 Analogs of the upper central series
Since $\lambda$ is a homomorphism on $(A,\circ)$, its kernel $\ker(\lambda)$ is a subgroup of $(A,\circ)$. But $\cdot$ and $\circ$ coincide on $\ker(\lambda)$. It then follows that $\ker(\lambda)$, and in particular $\mathrm{Soc}(A)$ in view of (4), is a subgroup of $(A,\cdot)$.
@@
    $\mathrm{Soc}(A)$ is normal in $(A,\circ)$: This follows from (6).
@@
We know that $\mathrm{Soc}(A)$ is a subgroup of $(A,\circ)$ by Proposition 4.1, so clearly $\mathrm{Ann}(A)$ is also a subgroup of $(A,\circ)$ by (5). Since $\cdot$ and $\circ$ coincide on $\ker(\lambda)$, it follows that $\mathrm{Ann}(A)$ is a subgroup of $(A,\cdot)$.
@@
$\mathrm{Soc}_{0}(A)=1,\quad\mathrm{Soc}(A/\mathrm{Soc}_{n}(A))=\mathrm{Soc}_{n+1}(A)/\mathrm{Soc}_{n}(A),$ (7)
@@
$\mathrm{Ann}_{0}(A)=1,\quad\mathrm{Ann}(A/\mathrm{Ann}_{n}(A))=\mathrm{Ann}_{n+1}(A)/\mathrm{Ann}_{n}(A).$ (8)
@@
$A_{1}=A,\quad A_{n+1}=A_{n}/\mathrm{Soc}(A_{n})$ (9)
instead, and the SocleSeries command in GAP [26] returns this series (9) rather than the series (7). But observe that by induction, we have a skew brace isomorphism, i.e. a map that is an automorphism with respect to both of the operations $\cdot$ and $\circ$, between
@@
Hence, the terms in the series (9) are nothing but quotients of those in the socle series of our definition, and there is no fundamental difference between the two series.
@@
$\mathrm{Ann}_{n}(A)\subseteq\mathrm{Soc}_{n}(A)$ (10)
@@
Let $A=(C_{p}\times C_{q},\cdot,\circ)$ be the brace in Example 3.1(i). From (4), we see that $\mathrm{Soc}(A)=\ker(\lambda)$ because $(A,\cdot)$ is abelian. Moreover, from (3), it is easy to see that $\mathrm{Soc}(A)=C_{p}\times\{1\}$ and the quotient $A/\mathrm{Soc}(A)$ is a trivial brace. This means that
@@
$\mathrm{Ann}_{n}(A)\subseteq\mathrm{Soc}_{n}(A)\subseteq\zeta_{n}(A,\cdot)\quad\mbox{and}\quad\mathrm{Ann}_{n}(A)\subseteq\zeta_{n}(A,\circ),$ (11)
@@ 5 Comparison of the different types of nilpotency
Nevertheless, there are some relations among the four different types of nilpotency. For example, from (10), it is clear that annihilator nilpotent implies socle nilpotent. Moreover, it was recently proven in [15] that a left nilpotent skew brace $A$ with nilpotent additive group $(A,\cdot)$ and $A^{3}=1$ is necessarily right nilpotent. But as we can see from Example 3.1, in general, there is no implication between left nilpotent and right nilpotent. Here, we shall survey a couple of known results about the various types of nilpotency.
@@ 2. $(b)$
    by (12), so it follows that $x\in\mathrm{Soc}_{c+j+1}(A)$, as desired.
@@
Now, assume that (b) holds. We have $\mathrm{Soc}_{n}(A)=A$ for some $n$ from the hypothesis. Note that $(A,\cdot)$ is nilpotent by (11). We also have
@@ 3. $(c)$
because $\mathrm{Soc}(A)$ is normal in $(A,\circ)$ by Proposition 4.1. On the other hand, recall from (6) that $a\circ x\circ\overline{a}=\lambda_{a}(x)$ and this lies in $\ker(\lambda)$. This yields
$[a,x]_{\circ}$ $=\lambda_{a}(x)\circ\overline{x}=\lambda_{a}(x)x^{-1}=a*x.$
@@
    by (13), which means that $x\in\mathrm{Ann}_{c+j+1}(A)$, as desired.
@@ 6 Relationship among the terms in the series
$a*y\in\mathrm{Ann}_{n}(A)*A^{n-(k+1)}\subseteq\mathrm{Ann}_{k+1}(A)$ (14)
@@
$=\varphi_{a}(\lambda_{x}(y))\varphi_{a}(y)^{-1}$ (15)
@@
and together with (14), this yields
@@
which can be assumed to be a homomorphism by induction on $k$ as in (14). In order to prove (F), we would have to show that
@@
as in (15) because $\lambda_{y}(x)$ and $x$ are not always elements of $A^{(n-(k+1))}$ here. There is no way to amend this problem. In fact, in general, (F) is false.
@@
$\mathrm{Im}(\psi_{b}-\mathrm{id}_{C})\subseteq\ker(\phi),$ (16)
@@
$(b,c)*(x,y)$ $=((\phi_{-c}-\mathrm{id}_{B})(x),(\psi_{b}-\mathrm{id}_{C})(y))$ (17)
$[(b,c),(x,y)]$ $=((\mathrm{id}_{B}-\phi_{y})(b)+(\phi_{c}-\mathrm{id}_{B})(x),0)$ (18)
@@
This proves the identities (17) and (18).
@@
with equalities when $b\in\langle\vec{e}_{3}\rangle$ but $b\neq\vec{0}$. We see that (16) is satisfied.
On the one hand, it is clear from (17) and the above that
@@
On the other hand, it is not hard to use (17) and (18) to verify that
$\mathrm{Ann}_{1}(A)$ $\supseteq\langle\vec{e}_{1}\rangle\times\langle\vec{e}_{1}\rangle,$ (19)
@@
The three inclusions in (19) follow from these observations.
Now, a simple calculation using (17) yields
@@ 7 Another lower central series and annihilator nilpotency
$\Gamma_{1}(A)=A,\quad\Gamma_{n+1}(A)=\langle\Gamma_{n}(A)*A,\,A*\Gamma_{n}(A),\,[A,\Gamma_{n}(A)]\rangle,$ (20)
@@
Therefore, perhaps (20) and the annihilator series (8) are, respectively, the “correct” analogs of the lower and upper central series, in the setting of skew braces. There is another reason to believe that this is indeed the case. A categorical approach was employed to study skew braces in [18], where it was shown that skew braces form a semi-abelian category, and the so-called Huq commutator of ideals is given as follows. Let us remark in passing that the categorical aspects of Hopf braces, an algebraic structure defined in [2] as a Hopf-theoretic generalization of skew braces, were also studied in [27].
@@
$\Gamma_{1}^{\prime}(A)=A,\quad\Gamma_{n+1}^{\prime}(A)=[A,\Gamma_{n}^{\prime}(A)]^{A},$ (21)
for all $n\in\mathbb{N}$. This coincides with the series (20). We note that Remark 7.6 and Proposition 7.7 were observed in [7] as well.
@@
Therefore, the series (20) can be defined using a commutator operator in the same way the lower central series of a group is defined. Our discussion so far also suggests that annihilator nilpotency is perhaps the more appropriate analog of nilpotency in the setting of skew braces.
@@
Let $I$ be an ideal of $A$. According to (21) and Proposition 7.7, as a skew brace, the lower central series of $I$ is defined by